OpenStudy (anonymous):

house blend coffee is 50% columbian beans and special blend coffee is 80% columbian beans. how much of each should be used to produce 100 kg of a blend that is 68% columbian beans

OpenStudy (anonymous):

Let house blend be H and special blend be S. eq 1 ----- H+S=100 eq 2 ----- as the percentages are 50 and 80 respectively, 50%H+80%S=68% 0.5H+0.8S=68 solve the equations simultaneously. eq 1 ------ S=100-H Substituting in eq 2 0.5H+0.8(100-H)=68 0.5H+80-0.8H=68 -0.3H=-12 H=40kg S=100-H=60kg